OpenStudy (jozelynw):

A tank currently holds 5 gallons of water. It is being filled at the rate of 1/4 gallon per minute . How long will it take for the tank to hold 9 gallons of water ?

OpenStudy (anonymous):

1) 5b + 6 = 21 5b = 15 b = 3 2) 9 - 5 = 4 gallons left to be filled .25x = 4 x = 16 minutes 3) x + 2/5 = 12 x = 12 - 2/5 = 60/5 - 2/5 = 58/5 4) 8 + a = 13 Subtract 8 from both sides (Both sides are equal, so performing the same operation on both sides is justified) 8 + a - 8 = 13 - 8 a = 5

OpenStudy (jozelynw):

How do you have the whole page?

OpenStudy (anonymous):

I've did this before

OpenStudy (jozelynw):

oh

OpenStudy (anonymous):

do u want me to give u a medal

OpenStudy (jozelynw):

What medal?

OpenStudy (anonymous):

just a medal

OpenStudy (anonymous):

for math

OpenStudy (jozelynw):

Sure , I guess.